# T-GVC: Trajectory-Guided Generative Video Coding at Ultra-Low Bitrates

### v2

@@ Abstract
Recent advances in video generation techniques have given rise to an emerging paradigm of generative video coding, aiming to achieve semantically accurate reconstructions in Ultra-Low Bitrate (ULB) scenarios by leveraging strong generative priors. However, most existing methods are limited by domain specificity (e.g., facial or human videos) or an excessive dependence on high-level text guidance, which often fails to capture motion details and results in unrealistic reconstructions. To address these challenges, we propose a Trajectory-Guided Generative Video Coding framework (dubbed T-GVC). T-GVC employs a semantic-aware sparse motion sampling pipeline to effectively bridge low-level motion tracking with high-level semantic understanding by extracting pixel-wise motion as sparse trajectory points based on their semantic importance, not only significantly reducing the bitrate but also preserving critical temporal semantic information. In addition, by incorporating trajectory-aligned loss constraints into diffusion processes, we introduce a training-free latent space guidance mechanism to ensure physically plausible motion patterns without sacrificing the inherent capabilities of generative models. Experimental results demonstrate that our framework outperforms both traditional codecs and state-of-the-art end-to-end video compression methods under ULB conditions. Furthermore, additional experiments confirm that our approach achieves more precise motion control than existing text-guided methods, paving the way for a novel direction of generative video coding guided by geometric motion modeling.
@@ 1 Introduction
The rapid advancement of video diffusion models (VDMs) (Xing et al. 2024b) has significantly intensified research interest in extending generative video coding to natural video content. Specifically, some recent works (Zhang et al. 2024b; Wan, Zheng, and Fan 2025) attempt to model spatio-temporal information solely based on texts and keyframes from the original video. However, although text serves as an effective modality for representing high-level semantics of video, text-guided generative models often encounter difficulties in reconstructing motion with high perceptual quality and may even produce visibly inferior results in certain cases. In parallel, image-to-video or image-text-to-video models (Xing et al. 2024a; Zhang et al. 2024a) have shown the potential to synthesize motion sequences conditioned on the structural and textual information derived from a pair of keyframes. Nevertheless, these methods still struggle to ensure temporal consistency and visual stability when applied to complex real-world scenarios. The generated outputs may exhibit unexpected artifacts, such as unnatural motion patterns or missing dynamic objects, compromising the temporal semantic fidelity with the original video. To allow for more precise generation, an intuitive idea is to take advantage of low-level motion information, as shown in Figure 1. Unfortunately, employing dense motion representation incurs a higher bitrate cost for VDMs to generate high-fidelity videos. It is essential to find an effective way to extract compact motion guidance which contains key temporal semantic information from the original video.
@@
In response to the critical challenge of generative video coding under ULB conditions, we propose a Trajectory-guided Generative Video Coding framework. To balance the trade-off between temporal semantic information preservation and bitrate saving, a semantic-aware sparse motion sampling pipeline is first proposed to bridge low-level motion tracking with high-level semantic understanding. Specifically, we track the motion of pixels on a pre-defined grid and classify them into distinct motion instances according to motion patterns. To further improve coding efficiency, a sparse sampling of motion instances is performed based on their semantic importance, representing motion as a set of trajectory points. The variations in trajectory points reflect different motion patterns (such as translation, occlusion, deformation, etc.), which preserves critical temporal semantic information of the original video while significantly reducing the bitrate. To guide the generative model in reconstructing videos aligned with target trajectories without compromising its inherent generative capabilities, we propose a training-free guidance approach. Unlike existing methods that constrain feature maps in intermediate network layers, we design a simple yet effective guidance function and directly impose constraints on the latent space to ensure that the trajectories of target points align with real-world motion paths.
@@ 3.3 Semantic-aware sparse motion sampling
Given two keyframes $\{K_{i},K_{i+1}\}$ and their intermediate frames, we design a bidirectional tracking scheme to extract dense motion trajectories with improved temporal consistency. For each clip $\{F_{0},F_{1},...,F_{L-1}\}$ with temporal length $L$, we initialize isometric grids (grid size $=N_{grid}\times N_{grid}$) in the first frame and employ Co-Tracker (Karaev et al. 2024) to trace pixel-wise displacements originating from the grid vertices. The generated forward trajectory sequences over $L$ frames are represented as temporal chains of 3D coordinate points:
@@ 3.4 Video reconstruction
Let $F\left(z_{t}^{i},\tau_{i}^{j}\right)$ denote the feature of the $i-th$ latent $z_{t}^{i}$ at time step t within the region covered by the sparse trajectories $\tau_{i}^{j}$ of the motion instance $\mathcal{T}^{j}$. To reconstruct the temporal semantic information in the guidance of trajectories, we define a loss function to measure how well the trajectories of $z_{t}$ align with the trajectory guidance, thereby constraining the denoising process:
@@ 4.2 Quantitative and qualitative results
We compare T-GVC with traditional codecs (H.265/HEVC (Sullivan et al. 2012) and H.266/VVC (Bross et al. 2021; Man et al. 2025)) and a neural video codec (DCVC-FM (Li, Li, and Lu 2024)) to evaluate its coding performance. The comparison results with other generative video coding schemes are provided and discussed in Section 4.3. For H.265, we use HM-18.0 with QP = 51, 45 and 39. For H.266, we use VTM-23.7 with QP = 63, 57, 51, 45 and 39. For DCVC-FM, we set the $q$ indexes to 0, 8, 16, 24 and 32, which control the compression level. All codecs above are tested under Low Delay P configuration with intra-period$=$-1.
The R-D curves for the test video classes are presented in Figure 5, where the bitrate of our T-GVC is controlled by adjusting the quality of keyframes. As observed, T-GVC outperforms both traditional codecs and the neural video codec in terms of perceptual reconstruction quality and semantic fidelity at ULB conditions. Figure 6 shows the visual results for the sequence ’$BasketballDrill\_832x480\_50$’ from the HEVC Class C dataset. To ensure a fair comparison, reconstructed sequences from different methods are selected at comparable bitrates. It is observed that the reconstructed videos generated by T-GVC demonstrate superior fidelity in the background regions and key foreground objects (e.g., basketball and the players in red). In contrast, VTM exhibits noticeable blocking artifacts while DCVC-FM suffers from significant high-frequency detail loss, leading to degraded texture quality. It is worth noting that although some texture and structural details in the reconstructed frames might be partially missing or inaccurate to some extent (e.g., the direction of the ball and the appearance of players in green), the overall visual quality remains satisfactory for the ULB scenarios.
@@ 4.3 Ablation study
To further verify the rationality of our design, we conduct ablation studies on two core components of the T-GVC framework: the guidance mechanism and the motion sampler module, as well as the keypoint quantity within each trajectory instance. Video sequences characterized by large motions, including $ParkScene$, $BasketballDrill$, $PartyScene$, $videoSRC02$ and $videoSRC05$ from HEVC Class B, Class C and MCL-JCV, are selected for evaluation. The coding efficiency is assessed using the Bjøntegaard Delta rate (BD-rate) (Bjontegaard 2001), where PSNR is replaced by $1-LPIPS$. Negative BD-rate values indicate bitrate saving for equivalent perceptual quality. We apply DynamiCrafter with empty prompts (i.e., only guided by keyframes) as the comparison anchor in our experiments. The reported BD-rate results are calculated by averaging all test sequences.
@@
In prior generative video coding methods (Zhang et al. 2024b; Wan, Zheng, and Fan 2025), text is commonly adopted as motion guidance. To demonstrate the superiority of our guidance mechanism, we implement three variants of generative frameworks based on DynamiCrafter: one guided by text (denoted as Text), one guided by trajectories (denoted as Trajectory) and the other by both text and trajectories (denoted as Text+Trajectory). Trajectories are sampled through our proposed sparse motion sampler, while text descriptions are generated from the source video using Hunyuan-Large (Sun et al. 2024) without limitation on length. The comparison results are presented in Table 2, where ”W/o motion bitrate” refer to excluding the bitrate consumed by motion-related information when calculating BPP, while ”W/ motion bitrate” indicate that this bitrate is included. It is observed that the trajectory-guided approach consistently achieves better coding performance under both evaluation conditions. In particular, the perceptual quality improvement brought by trajectory guidance makes up for the additional bitrate overhead, offering a distinct advantage over text-based guidance. When overhead is not considered, the trajectory-guided method still outperforms the text-guided method. Combining text and trajectory guidance may enhance texture on simple-motion sequences. Some visual examples in Figure 7 further demonstrate that trajectory guidance can perform more precise motion control, especially in dynamic areas. Text guidance can improve perceptual quality to some extent and combining text and trajectory guidance may enhance texture on simple-motion sequences. However, in more complex cases, text prompts may introduce artifacts that conflict with trajectory-aligned motion, leading to degraded results. For instance, as shown in the second column of Figure 7, the basketball appears behind the net, which violates the physical plausibility.

### v3

@@ Abstract
Recent advances in video generation techniques have given rise to an emerging paradigm of generative video coding for Ultra-Low Bitrate (ULB) scenarios by leveraging powerful generative priors. However, most existing methods are limited by domain specificity (e.g., facial or human videos) or excessive dependence on high-level text guidance, which tend to inadequately capture fine-grained motion details, leading to unrealistic or incoherent reconstructions. To address these challenges, we propose Trajectory-Guided Generative Video Coding (dubbed T-GVC), a novel framework that bridges low-level motion tracking with high-level semantic understanding. T-GVC features a semantic-aware sparse motion sampling pipeline that extracts pixel-wise motion as sparse trajectory points based on their semantic importance, significantly reducing the bitrate while preserving critical temporal semantic information. In addition, by integrating trajectory-aligned loss constraints into diffusion processes, we introduce a training-free guidance mechanism in latent space to ensure physically plausible motion patterns without sacrificing the inherent capabilities of generative models. Experimental results demonstrate that T-GVC outperforms both traditional and neural video codecs under ULB conditions. Furthermore, additional experiments confirm that our framework achieves more precise motion control than existing text-guided methods, paving the way for a novel direction of generative video coding guided by geometric motion modeling.
@@ 1 Introduction
The rapid advancement of video diffusion models (VDMs) (Xing et al. 2024b) has significantly intensified research interest in extending generative video coding to natural video content. Specifically, some recent works (Zhang et al. 2024b; Wan, Zheng, and Fan 2025) attempt to model spatio-temporal information solely based on texts and keyframes from the original video. However, although text serves as an effective modality for representing high-level semantics of video, text-guided generative models often encounter difficulties in reconstructing motion with high perceptual quality and may even produce visibly inferior results in certain cases. In parallel, image-to-video or image-text-to-video models (Xing et al. 2024a; Zhang et al. 2024a) have shown the potential to synthesize motion sequences conditioned on the structural and textual information derived from a pair of keyframes. Nevertheless, these methods still struggle to ensure temporal consistency and visual stability when applied to complex real-world scenarios. The generated outputs may exhibit unexpected artifacts, such as unnatural motion patterns or missing dynamic objects, compromising the temporal semantic fidelity with the original video. To allow for more precise generation, an intuitive idea is to take advantage of low-level motion information, as shown in Figure 1. Unfortunately, employing dense motion representation incurs a higher bitrate cost for VDMs to generate high-fidelity videos. It is crucial to develop an effective strategy to extract compact motion guidance that preserves key temporal semantic information from the original video.
@@
In response to the critical challenge of generative video coding under ULB conditions, we propose a Trajectory-guided Generative Video Coding framework. To balance the trade-off between temporal semantic information preservation and bitrate saving, a semantic-aware sparse motion sampling pipeline is first proposed to bridge low-level motion tracking with high-level semantic understanding. Specifically, we track the motion of pixels on a pre-defined grid and classify them into distinct motion instances according to motion patterns. To further improve coding efficiency, a sparse sampling of motion instances is performed based on their semantic importance, representing motion as a set of trajectory points. The variations in trajectory points reflect different motion patterns (such as translation, occlusion, deformation, etc.), which preserves critical temporal semantic information of the original video while significantly reducing the bitrate. To guide the diffusion-based generative model in reconstructing motion-aligned videos without compromising its inherent generative capabilities, we propose a training-free guidance approach. In contrast to existing methods that constrain on intermediate feature maps, our approach directly imposes constraints on the latent space of diffusion model via a lightweight yet effective guidance function, ensuring that the overall trajectories of target motion instances align with real-world motion paths.
@@ 3.3 Semantic-aware sparse motion sampling
Given two keyframes $\{K_{i},K_{i+1}\}$ and their intermediate frames, we design a bidirectional tracking scheme to extract dense motion trajectories with improved temporal consistency. For each clip $\{F_{0},F_{1},...,F_{L-1}\}$ with temporal length $L$, we initialize isometric grids (grid size $N_{grid}\times N_{grid}$) in the first frame and employ Co-Tracker (Karaev et al. 2024) to trace pixel-wise displacements originating from the grid vertices. The generated forward trajectory sequences over $L$ frames are represented as temporal chains of 3D coordinate points:
@@ 3.4 Video reconstruction
Let $F\left(z_{t}^{i},\tau_{i}^{j}\right)$ denote the feature of the ithi-th latent $z_{t}^{i}$ at time step t within the region covered by the sparse trajectories $\tau_{i}^{j}$ of the motion instance $\mathcal{T}^{j}$. To reconstruct the temporal semantic information in the guidance of trajectories, we define a loss function to measure how well the trajectories of $z_{t}$ align with the trajectory guidance, thereby constraining the denoising process:
@@ 4.2 Quantitative and qualitative results
We compare T-GVC with traditional codecs (H.265 (Sullivan et al. 2012) and H.266 (Bross et al. 2021)) and one of the most competitive open-source neural video codec (DCVC-FM (Li, Li, and Lu 2024)) with a wide bitrate range to evaluate its coding performance at ULB. The comparison results with other generative video coding schemes are provided and discussed in Section 4.3. For H.265, we use HM-18.0 with QP = 51, 45 and 39. For H.266, we use VTM-23.7 with QP = 63, 57, 51, 45 and 39. For DCVC-FM, we set the $q$ indexes to 0, 8, 16, 24 and 32, which control the compression level. All codecs above are tested under Low Delay P configuration with intra-period==-1.
The R-D curves for the test video classes are presented in Figure 5, where the bitrate of our T-GVC is controlled by adjusting the quality of keyframes. As observed, T-GVC outperforms both traditional codecs and the neural video codec in terms of perceptual reconstruction quality and semantic fidelity at ULB conditions. Moreover, T-GVC achieves lower bitrate points (below 0.005 bpp) than GAN- or VQ-VAE-based methods (Yang, Timofte, and Van Gool 2022; Qi et al. 2025), with detailed comparisons provided in the appendix. Figure 6 shows the visual results for the sequence ‘BasketballDrill_832x480_50’ from the HEVC Class C dataset. To ensure a fair comparison, reconstructed sequences from different methods are selected at comparable bitrates. It is observed that the reconstructed videos generated by T-GVC demonstrate superior fidelity in the background regions and key foreground objects (e.g., basketball and the players in red). In contrast, VTM exhibits noticeable blocking artifacts while DCVC-FM suffers from significant high-frequency detail loss, leading to degraded texture quality. It is worth noting that although some texture and structural details in the reconstructed frames might be partially missing or inaccurate to some extent (e.g., the direction of the ball and the appearance of players in green), the overall visual quality remains satisfactory for the ULB scenarios.
@@ 4.3 Ablation study
To further verify the rationality of our design, we conduct ablation studies on two core components of the T-GVC framework: the guidance mechanism and the motion sampler module, as well as the keypoint quantity within each trajectory instance. Video sequences characterized by large motions, including $ParkScene$, BasketballDrillBasketballDrill, PartyScenePartyScene, $videoSRC02$ and $videoSRC05$ from HEVC Class B, Class C and MCL-JCV, are selected for evaluation. The coding efficiency is assessed using the Bjøntegaard Delta rate (BD-rate) (Bjontegaard 2001), where PSNR is replaced by 1LPIPS1-LPIPS. Negative BD-rate values indicate bitrate saving for equivalent perceptual quality. We apply DynamiCrafter with empty prompts (i.e., only guided by keyframes) as the comparison anchor in our experiments. The reported BD-rate results are calculated by averaging all test sequences.
@@
In prior generative video coding methods (Zhang et al. 2024b; Wan, Zheng, and Fan 2025), text is commonly adopted as motion guidance. To demonstrate the superiority of our guidance mechanism, we implement three variants of generative frameworks based on DynamiCrafter: one guided by text (denoted as Text), one guided by trajectories (denoted as Trajectory) and the other by both text and trajectories (denoted as Text+Trajectory). Trajectories are sampled through our proposed sparse motion sampler, while text descriptions are generated from the source video using Hunyuan-Large (Sun et al. 2024) without limitation on length. The comparison results are presented in Table 2, where ”W/o motion bitrate” refer to excluding the bitrate consumed by motion-related information when calculating BPP, while ”W/ motion bitrate” indicate that this bitrate is included. It is observed that the trajectory-guided approach consistently achieves better coding performance under both evaluation conditions. In particular, the perceptual quality improvement brought by trajectory guidance makes up for the additional bitrate overhead, offering a distinct advantage over text-based guidance. When overhead is not considered, the trajectory-guided method still outperforms the text-guided method. Combining text and trajectory guidance may enhance texture on simple-motion sequences. Some visual examples in Figure 7 further demonstrate that trajectory guidance can perform more precise motion control, especially in dynamic areas. Text guidance can improve perceptual quality to some extent and combining text and trajectory guidance may enhance texture on simple-motion sequences. However, in more complex cases, text prompts may introduce artifacts that conflict with trajectory-aligned motion, leading to degraded results. For instance, as shown in the first column of Figure 7, the basketball appears behind the net, which violates the physical plausibility.